Find the critical value (or values) for the 𝑑-test for each.


a. 𝑛 = 10 ; ∝= 0.05 ; right tailed

b. 𝑛 = 18 ; ∝= 0.10 ; two-tailed

c. 𝑛 = 6 ; ∝= 0.01 ; left-tailed

d. 𝑛 = 9 ; ∝= 0.025 ; right tailed

e. 𝑛 = 15 ; ∝= 0.05 ; two-tailedΒ 


1
Expert's answer
2021-10-05T15:38:10-0400

a. "\ud835\udc5b = 10,\u221d= 0.05 , df=n-1=10-1=9," right tailed

"t_c=1.833113"

b. "\ud835\udc5b = 18,\u221d= 0.10 , df=n-1=18-1=17," two tailed

"t_c=1.739607"

c. "\ud835\udc5b = 6,\u221d= 0.01 , df=n-1=6-1=5," left tailed

"t_c=-3.364762"


d. "\ud835\udc5b = 9,\u221d= 0.025 , df=n-1=9-1=8," right tailed

"t_c=2.306002"


e. "\ud835\udc5b = 15,\u221d= 0.05 , df=n-1=15-1=14," two tailed

"t_c=2.144787"